# Nonparametric Jackknife Instrumental Variable Estimation and Confounding Robust Surrogate Indices

### v1

@@ 5 Semiparametric cross-fold estimator
$\widehat{\theta}=\frac{1}{n^{\prime}}\sum_{i=1}^{n^{\prime}}\widehat{h}(S_{i}^{(K+1)})+\frac{4}{N}\sum_{i}\xi(S_{i})(Y_{i}^{\prime}-\widehat{h}(S_{i}^{\prime})).$ (18)
@@
Suppose that $\|\Pi(\widehat{h}-h_{0})\|\|\Pi(\widehat{\xi}-\xi_{0})\|=o_{P}((n^{\prime}\wedge N)^{-1/2})$ and that $\|\widehat{h}-h_{0}\|=o_{P}(1)$ and $\|\widehat{\xi}-\xi_{0}\|=o_{P}(1)$ Then
@@
where $\sigma_{1}^{2}=\mathrm{Var}(h_{0}(S)\mid A=a^{\prime})$ and $\sigma_{2}^{2}=\mathrm{Var}(\xi_{0}(S)\mid A\in[K])\times\mathrm{Var}(Y-h_{0}(S))\mid A\in[K])$.
@@ 6 Numerical Experiments
We perform simulation experiments to study the behavior of cross-fold npJIVE. To begin, we fix a data-generating process with treatment effects $\mu_{a}$ generated from a normal distribution with variance $\sigma_{a}$, where the latter choice controls the strength of the instruments. The short-term outcome $S$ is discrete and generated from a multinomial distribution with probability vector drawn $\text{Dirichlet}(\mathbf{1}\cdot\exp(\mu_{a})$. The long-term outcome $Y$ is generated as a combination of non-linear function $h(S)=S+5\times(S>3)$, an unobserved confounder $U\sim\mathcal{N}(0,1)$, and standard normal noise. In the ‘historical experiments’ dataset $\mathfrak{H}$, we have $A=[1,\dots k]$ experiments that contain $n=50$ observations each that follows the above DGP. In a ‘target experiments’ dataset $\mathfrak{T}$, we have $n=100$ observations from the same DGP.
To begin, we compare the performance of a kernel-ridge implementation of $\hat{h}(S)$ trained on the historical data $\mathfrak{H}$ in recovering the true $h_{0}(S)$ in the target data $\mathfrak{T}$ using the weak norm $\frac{1}{K}\sum_{a}\hat{\mathbb{E}}(Y-\hat{h}(S)\mid A=a]^{2}/\sigma_{a}^{2}$ varying instrument strength $\sigma_{a}$ and number of instruments $k$. We report this weak norm in figure 2, and find that the cross-fold moments estimator empirically converges in the weak norm, with faster convergence when instruments are stronger.
Figure 2: Weak norm normalized by instrument strength, that is $\frac{1}{K}\sum_{a}\hat{\mathbb{E}}(Y-\hat{h}(S)\mid A=a]^{2}/\sigma_{a}^{2}$ varying instrument strength $\sigma_{a}$ and number of instruments evaluated on target experiments
Next, we perform a simulation experiments to compare the performance of NPJIVE with its non-cross-fold counterpart, the sieve minimum distance (SMD) estimator (Ai and Chen, 2003). Both estimators of $\hat{h}(S)$ involve fitting a kernel-ridge regression of the form $\hat{\beta}_{h}=(\Phi Q\Phi+\lambda I_{n})^{-1}\Phi QY$ on the historical experiments data, where $\Phi$ is the Gram matrix of the short-term outcomes $S$, and $Q$ is sparse selector matrix of dimensions $nK\times nK$ (where $n$ is the sample size in each experiment and $K$ is the number of instruments). The sparsity of $Q$ distinguishes our estimator from 2SLS/SMD: without cross-folds, $Q$ is block-diagonal with $n\times n$ matrices of ones on the diagonal, while the use of cross-folds introduces zeros to compute held-out averages (with maximal sparsity with two folds and minimal with $n-1$ folds, which is the conventional leave-one-out jackknife), as illustrated in fig 4.
Similar to above, we fix a DGP with treatment effects $\mu_{a}$ drawn from a normal distribution with variance $\sigma_{a}$, which governs instrument strength, and the short term outcome is generated from a uniform distribution $\mathsf{Unif}[\mu_{a}-0.5,\mu_{a}+0.5]$, and the long-term outcome is generated as a combination of a non-linear function $h(S)=S+0.5(S>0.25)-\sin(S)$ and an unobserved confounder $U$. In the ‘historical experiments’ dataset $\mathfrak{H}$, we have $A=[1,\dots,K=5000]$ experiments, each with $n$ observations each, where we shall vary $n$ to ‘very small’ and ‘moderate’ values. We compare the performance of SMD and npJIVE in recovering $\widehat{S}$ in the weak norm. We report results for two levels of instrument strength $\sigma_{a}$, and experiment size $n$, in figure 4, where the x-axis reports the bandwidth of the kernel (with the optimal value denoted by a dotted line). We find that a tuned npJIVE consistently weakly outperforms SMD, with large gaps with small weak experiments.
Figure 3: Q matrices for KRR implementation of SMD and npJIVE. Each matrix is $nK\times nK$, with black cells corresponding with zeroes and light cells corresponding with ones.
l Refer to caption
Figure 4: Weak norm of npJIVE and SMD, varying RBF kernel (inverse)bandwidth $\gamma$, instrument strength $\sigma_{a}$ (column) and sample size in each experiment (rows), with total number of instruments fixed at $5000$. We see that a tuned (i.e. optimal bandwidth) npJIVE consistently outperforms SMD, with gaps being largest when instruments are weak and experiments are small
7 Concluding remarks
The proliferation of experimentation on digital platforms provides a bountiful source of instruments (Peysakhovich and Eckles, 2018). This holds great potential to learn much more than what each such experiment taught us about the treatment that was experimented with. In particular, using it to understand the relationship between short- and long-term metrics (Bibaut et al., 2024), we may even be able to predict the effects of new experiments before observing any long-term outcome, even when the perfect surrogacy needed for observational surrogate index methods (Athey et al., 2019; Chen and Ritzwoller, 2023) fails. A key challenge is that, while we have many experiments, each experiment only tells us so little. Here we tackled the combined challenges of NPIV, many weak IVs, and inference on low dimensional functionals, our solution thereto newly enables statistical inference on long-term causal effects with nonparametric surrogate indices learned from many weak historical experiments.
[…]
@@ Appendix A Proof of theorem 4
That $\overline{\mathcal{R}(\Pi^{*})}=\mathcal{R}(\Pi^{*})$ directly follows from the fact that $\mathcal{R}(\Pi^{*})$ is a finite dimensional linear subspace, and therefore equals its closure.
Let $\phi_{a}:s\mapsto p[s\mid a]$. Let $r=\mathrm{rank}(\{\phi_{a}:a\in[K]\})$. There exists $a_{1},\ldots,a_{r}$ such that $\phi_{a_{1}},\ldots,\phi_{a_{r}}$ are linearly independent.
From now on, we proceed in four steps.
Step 1: $\Pi\phi_{a_{1}},\ldots,\Pi\phi_{a_{r}}$ are linearly independent.
Let $\gamma_{1},\ldots,\gamma_{r}$ be such that $\sum_{i=1}^{r}\gamma_{i}\Pi\phi_{a_{i}}=0$. In particular must have that, for every $j=1,\ldots,r$, $\sum_{i=1}^{r}\gamma_{i}\Pi\phi_{a_{i}}(a_{j})=0$, and therefore, by summation, $\sum_{1\leq i,j\leq r}\gamma_{i}\gamma_{j}\Pi\phi_{a_{i}}(a_{j})=0$, that is, by definition of $\Pi$,
$\sum_{1\leq i,j\leq r}\gamma_{i}\gamma_{j}\langle\phi_{a_{i}},\phi_{a_{j}}\rangle_{L_{2}(\mathcal{S})}=\langle\sum_{i=1}^{r}\gamma_{i}\phi_{a_{i}},\sum_{i=1}^{r}\gamma_{i}\phi_{a_{i}}\rangle_{L_{2}(\mathcal{S})}=0.$ (20)
Therefore, from positive definiteness, $\sum_{i=1}^{r}\gamma_{i}\phi_{a_{i}}$, which implies that $\gamma_{1}=\ldots=\gamma_{r}=0$, since $\phi_{a_{1}},\ldots,\phi_{a_{r}}$ are linearly independent. We have thus shown that $\Pi\phi_{a_{1}},\ldots,\Pi\phi_{a_{r}}$ are linearly independent.
Step 2: $\Pi^{*}\Pi\phi_{a_{1}},\ldots,\Pi^{*}\Pi\phi_{a_{r}}$ are linearly independent.
Let $\gamma_{1},\ldots,\gamma_{r}$ be such that $\sum_{i=1}^{r}\gamma_{i}\Pi^{*}\Pi\phi_{a_{i}}=0$. We must then have for every $j=1,\ldots,r,$ that
$\langle\sum_{i=1}^{r}\gamma_{i}\Pi^{*}\Pi\phi_{a_{i}},\phi_{a_{j}}\rangle_{L_{2}(\mathcal{S})}=0,$ (21)
and therefore, by summation, that
$\langle\sum_{i=1}^{r}\gamma_{i}\Pi^{*}\Pi\phi_{a_{i}},\sum_{i=1}^{r}\gamma_{i}\phi_{a_{j}}\rangle_{L_{2}(\mathcal{S})}=0,$ (22)
and thus, by definition of the adjoint, that
$\langle\sum_{i=1}^{r}\gamma_{i}\Pi\phi_{a_{i}},\sum_{i=1}^{r}\gamma_{i}\Pi\phi_{a_{i}}\rangle=0.$ (23)
By positive definiteness, $\sum_{i=1}^{r}\gamma_{i}\Pi\phi_{a_{i}}=0$, which implies that $\gamma_{1}=\ldots=\gamma_{r}=0$ since $\Pi\phi_{a_{1}},\ldots,\Pi\phi_{a_{r}}$ are linearly independent.
Step 3: $\mathrm{rank}(\Pi^{*})\leq r$.
We have that $\mathcal{R}(\Pi^{*})=\mathrm{Span}(s\mapsto p[a\mid s]:a\in[K])$. From Bayes formula,
$\mathcal{R}(\Pi^{*})=\mathrm{Span}\left(s\mapsto\frac{\phi_{a}(s)}{Kp(s)}:a\in[K]\right)\subseteq\mathrm{Span}\left(s\mapsto\frac{\phi_{a_{i}}(s)}{Kp(s)}:i=1,\ldots,r\right),$ (24)
since $\phi_{a_{1}},\ldots,\phi_{a_{r}}$ is a basis of $\mathrm{Span}(\phi_{a}:a\in[K])$.
Concluding.
We have that $\mathcal{R}(\Pi^{*}\Pi)\subseteq\mathcal{R}(\Pi^{*})$. From step 2, $\mathrm{rank}(\Pi^{*}Pi)\geq r$, and from step 3, $\mathrm{rank}(\Pi^{*})\leq r$. Therefore, we must have $\mathcal{R}(\Pi^{*}\Pi)=\mathcal{R}(\Pi^{*})$, which concludes the proof.
@@ B.2 Decomposition of the difference between population and empirical risk
$=$ $\langle\widehat{\Pi}_{0}(h_{0}-h+\eta),\widehat{\Pi}_{1}(h_{0}-h+\eta)\rangle-\langle\Pi(h_{0}-h),\Pi(h_{0}-h)\rangle$ (34)
$=$ $\langle(\widehat{\Pi}_{0}-\Pi)(h-h_{0}),(\widehat{\Pi}_{1}-\Pi)(h-h_{0})\rangle$ (35)
$+\sum_{v=0,1}\langle\Pi(h-h_{0}),(\widehat{\Pi}_{v}-\Pi)(h-h_{0})\rangle$ (36)
$+\sum_{v=0,1}\langle\Pi(h-h_{0}),(\widehat{\Pi}_{v}-\Pi)\eta\rangle$ (37)
$+\sum_{v=0,1}\langle(\widehat{\Pi}_{1-v}-\Pi)(h-h_{0}),(\widehat{\Pi}_{v}-\Pi)\eta\rangle$ (38)
$+\langle(\widehat{\Pi}_{0}-\Pi)\eta,(\widehat{\Pi}_{1}-\Pi)\eta\rangle$ (39)
@@ B.3 U-process terms
Denote $\mathcal{G}=\mathcal{H}-h_{0}$. For any $g\in\mathcal{G}$, let $g^{0}=g-\Pi g$, and let $\mathcal{G}^{0}=\{g-\Pi g:g\in\mathcal{G}\}$ For any such $g^{0}$, let
$Z_{N,1}^{0}(g)=\langle\widehat{\Pi}_{0}g^{0},\widehat{\Pi}_{1}g^{0}\rangle.$ (40)
@@
$[\widehat{\Pi}_{0}g^{0}](a)=O(\widetilde{\delta}_{n}\|g^{0}\|_{2,a}+\widetilde{\delta}_{n}^{2}).$ (43)
@@
$Z_{N,1}(g^{0})\leq\widehat{\delta}_{N}\|\widehat{\Pi}_{0}g^{0}\times g\|_{|\mathcal{D}_{0}}+\widehat{\delta}_{N}^{2}.$ (44)
@@
$\|\widehat{\Pi}_{0}g^{0}\times g\|_{|\mathcal{D}_{0}}=$ $\sqrt{\frac{2}{N}\sum_{i:V_{i}=1}\left\lvert\widehat{\Pi}_{0}g^{0}(A_{i})\right\rvert^{2}\mathbb{E}[(g^{0}(S_{i}))^{2}]}$ (45)
@@
$Z_{N,2}^{0}(g)=\langle\widehat{\Pi}_{0}\eta,\widehat{\Pi}_{1}g^{0}\rangle$ (49)
@@
$\max_{a\in[K]}|[\widehat{\Pi}_{0}\eta](a)|=O\left(\sqrt{\frac{\log K}{n}}\right).$ (51)
@@
$Z_{N,2}^{0}(g)=\frac{1}{N/2}\sum_{i:V_{i}=1}[\widehat{\Pi}_{0}\eta](A_{i})g^{0}(S_{i}).$ (52)
The conclusion then follows by applying lemma 8 conditional on $\mathcal{D}_{0}$, on the event that the above display holds, with $b_{i}=[\widehat{\Pi}_{0}\eta](A_{i})|$. ∎

### v2

@@ 5 Semiparametric cross-fold estimator
$\widehat{\theta}=\frac{1}{n^{\prime}}\sum_{i=1}^{n^{\prime}}\widehat{h}(S_{i}^{\mathrm{new}})+\frac{4}{N}\sum_{i}\xi(S_{i})(Y_{i}^{\prime}-\widehat{h}(S_{i}^{\prime})).$ (28)
@@
Suppose that $\|\mathcal{T}_{K}(\widehat{h}-h_{0})\|\|\mathcal{T}_{K}(\widehat{\xi}-\xi_{0})\|=o_{P}((n^{\prime}\wedge N)^{-1/2})$ and that $\|\widehat{h}-h_{0}\|=o_{P}(1)$ and $\|\widehat{\xi}-\xi_{0}\|=o_{P}(1)$ Then
@@
where $\sigma_{1}^{2}=\mathrm{Var}(h_{0}(S)\mid A=a^{\prime})$ and $\sigma_{2}^{2}=\mathrm{Var}(\xi_{0}(S)\mid A)\times\mathrm{Var}(Y-h_{0}(S))$.
@@ 6 Numerical Experiments
In this section, we consider the confounded surrogate setting of section 3. We demonstrate inconsistency of minimum distance estimators and consistency of npJIVE under many-weak-IVs asymptotics, and and debiasing via the one-step estimator.
We consider a one dimensional vector of short-term variables $S$. We let the structural parameter $h^{\star}$ be the function $h^{\star}:s\mapsto s+\sin(s)+\bm{1}\{s>0.25\}$. For any given $K$, we sample the first-stage effects $\Gamma_{1}$,…,$\Gamma_{K}$ on $S$ as i.i.d. draws from a normal $\mathcal{N}(0,\sigma_{\Gamma}^{2})$. For $i=1,\ldots,N$, we let $A_{i}=\left\lfloor i/n\right\rfloor$. We draw $S_{i}$ and $Y_{i}$ according to the following set of structural equations:
$U_{i}\sim$ $\mathcal{N}(0,\sigma_{U}^{2}),$ (30)
$S_{i}\sim$ $\mathrm{Uniform}(\Gamma_{A_{i}}-0.5,\Gamma_{A_{i}}+0.5)+U_{i},$ (31)
$Y_{i}=$ $h(S_{i})-U_{i}.$ (32)
The random variable $U_{i}$ is an unmeasured confounder. The draws $U_{1},\ldots,U_{N}$ are independent. We generate independent draws $S_{1}^{\mathrm{new}},\ldots,S^{\mathrm{new}}_{n^{\mathrm{new}}}$ according to
$U_{i}^{\mathrm{new}}\sim$ $\mathcal{N}(0,\sigma_{U}^{2}),$ (33)
$S_{i}^{\mathrm{new}}\sim$ $\mathrm{Uniform}(\Gamma^{\mathrm{new}}-0.5,\Gamma^{\mathrm{new}}+0.5)+U_{i}$ (34)
We take $\mathcal{H}$ to be the reproducing kernel Hilbert space (RKHS) induced by the Gaussian kernel with bandwidth $\nu>0$.
In all the experiments below, we set $\sigma_{\Gamma}=2$, $\Gamma^{\mathrm{new}}=1$, $\sigma_{U}=1$.
Figure 3: $n=30$, $\lambda_{h}=10^{-2}$, $\nu_{h}=1/3$, $L_{h}=5$, $\nu_{\mathrm{IPW}}=1/10$, $L_{\mathrm{IPW}}=10$, $w=3$
Figure 4: $n=100$, $\lambda_{h}=10^{-2}$, $\nu_{h}=1/4$, $L_{h}=7$, $\nu_{\mathrm{IPW}}=1/10$, $L_{\mathrm{IPW}}=10$, $w=3$
Figure 5: $n=300$, $\lambda_{h}=10^{-1}$, $\nu_{h}=1/10$, $L_{h}=10$, $\nu_{\mathrm{IPW}}=1/10$, $L_{\mathrm{IPW}}=10$, $w=3$
Figure 6: $n=3000$, $\lambda_{h}=10^{-1}$, $\nu_{h}=1/10$, $L_{h}=10$, $\nu_{\mathrm{IPW}}=1/10$, $L_{\mathrm{IPW}}=10$, $w=3$
As can be observed in the above figures, while they appear to be consistent as $n\to\infty$ and to significantly reduce median squared error when used to construct the one-step, the approximate-identification debiasing nuisances do not appear to be consistent as $K\to\infty$. This is in contrast with the the exact-identification debiasing nuisance which, under exact identification is actually consistent, as guaranteed by the results of section 4. We demonstrate empirically the consistency of the exact identification debiasing nuisance in a numerical experiment in which we ensure exact identification. For each cell $a$, we draw a length-5 vector of fist-stage effects
$\mu_{a}\sim\mathrm{Dirichlet}(10\times(1,\ldots,1)).$ (35)
Let $s_{m}=-1+2m/5$ for $m=1,\ldots,5$. Conditional on the first-stage effects, the data-generating process is as follows:
$U_{i}$ $\sim\mathrm{Unif}(-0.2,0.2)$ (36)
$S_{i}$ $\sim s_{\mathrm{Multinomial}(\mu_{a})}+U_{i}$ (37)
$Y_{i}$ $=h(S_{i})-10U_{i}.$ (38)
We generate the data in the target novel experiment cell according to
$U_{i}^{\mathrm{new}}$ $\sim\mathrm{Unif}(-0.2,0.2)$ (39)
$S_{i}^{\mathrm{new}}$ $\sim s_{\mathrm{Multinomial}(\mu^{\mathrm{new}})}+U_{i}^{\mathrm{new}},$ (40)
where $\mu^{\mathrm{new}}=(0,1/10,2/10,3/10,4/10)$. We report mean squared error, squared bias and variance as a function of $K$, under $n=100$ for this data-generating distribution in figure 7.
Figure 7: Exact identification.
@@ Appendix A Proof of theorem 4
That $\overline{\mathcal{R}(\mathcal{T}_{K}^{*})}=\mathcal{R}(\mathcal{T}_{K}^{*})$ directly follows from the fact that $\mathcal{R}(\mathcal{T}_{K}^{*})$ is a finite dimensional linear subspace, and therefore equals its closure.
For every $a\in[K]$, let $q_{a}=\bm{1}\{\cdot=a\}$, and let $q_{a}^{*}=\mathcal{T}_{K}^{*}q_{a}$. Let $r$ be the rank of $q_{1}^{*},\ldots,q_{K}^{*}$. Since $q_{1},\ldots,q_{K}$ form a basis of $\mathcal{Q}$, $q_{1}^{*},\ldots,q_{K}^{*}$ form a basis of $\mathcal{R}(\mathcal{T}_{K}^{*})$ and therefore $\mathcal{R}(\mathcal{T}_{K}^{*})$ has dimension $r$. Since $\mathcal{R}(\mathcal{T}_{K}^{*}\mathcal{T}_{K})\subseteq\mathcal{R}(\mathcal{T}_{K}^{*})$, the claim will follow if we can show that $\mathcal{R}(\mathcal{T}_{K}^{*}\mathcal{T}_{K})$ has dimension at least $r$ too.
Without loss of generality, suppose that $q_{1}^{*},\ldots,q_{r}^{*}$ are linearly independent. Let us show that $\mathcal{T}_{K}q_{1}^{*},\ldots,\mathcal{T}_{K}q_{r}^{*}$ are linearly independent too. Observe that for any $a,a^{\prime}$,
$\mathcal{T}_{K}q_{a}^{*}(a^{\prime})=K\langle\mathcal{T}_{K}q_{a}^{*},q_{a}\rangle_{L_{2}(\mathcal{Q})}=K\langle q_{a}^{*},q_{a^{\prime}}^{*}\rangle_{L_{2}(\mathcal{S})}.$ (41)
Therefore, for any $\beta_{1},\ldots,\beta_{r}\in\mathbb{R}$ such that $\sum_{a=1}^{r}\beta_{a}\mathcal{T}_{K}q_{a}^{*}=0$, we have
$0=\sum_{a^{\prime}=1}^{r}\beta_{a^{\prime}}\sum_{a=1}^{r}\beta_{a}\mathcal{T}_{K}q_{a}^{*}(a^{\prime})=K\langle\sum_{a=1}^{r}\beta_{a}q_{a}^{*},\sum_{a=1}^{r}\beta_{a}q_{a}^{*}\rangle_{L_{2}(\mathcal{S})},$ (42)
which implies that $\sum_{a=1}^{r}\beta_{a}q_{a}^{*}=0$, which in turns implies that $\beta_{1}=\ldots=\beta_{r}=0$ as $q_{1}^{*},\ldots,q_{r}^{*}$ are linearly independent, therefore $\mathcal{T}_{K}q_{1}^{*},\ldots,\mathcal{T}_{K}q_{r}^{*}$ are linearly independent. We have thus shown that application of $\mathcal{T}_{K}$ to $q_{1}^{*},\ldots,q_{r}^{*}$ didn’t reduce the rank.
It remains to show that application of $\mathcal{T}_{K}^{*}$ to $\mathcal{T}_{K}q_{1}^{*},\ldots,\mathcal{T}_{K}q_{r}^{*}$ doesn’t reduce the rank either. Suppose that $\beta_{1},\ldots,\beta_{r}$ are such that $\sum_{a=1}^{r}\beta_{a}\mathcal{T}_{K}^{*}\mathcal{T}_{K}q_{a}^{*}=0$. Taking inner products against $q_{1}^{*},\ldots,q_{r}^{*}$, we have
$0=\sum_{a^{\prime}=1}^{r}\beta_{a^{\prime}}\langle\sum_{a=1}^{r}\beta_{a}\mathcal{T}_{K}^{*}\mathcal{T}_{K}q_{a}^{*},q_{a^{\prime}}^{*}\rangle=\langle\sum_{a=1}^{r}\beta_{a}\mathcal{T}_{K}q_{a}^{*},\sum_{a=1}^{r}\beta_{a}\mathcal{T}_{K}q_{a}^{*}\rangle,$ (43)
which implies $\sum_{a=1}^{r}\beta_{a}\mathcal{T}_{K}q_{a}^{*}=0$ and in turns implies $\beta_{1}=\ldots=\beta_{r}=0$ since $\mathcal{T}_{K}q_{1}^{*},\ldots\mathcal{T}_{K}q_{r}^{*}$ are linearly independent. Therefore, $\mathcal{R}(\mathcal{T}_{K}^{*}\mathcal{T}_{K})$ must have dimension at least $r$ which concludes the proof. ∎
@@ B.2 Decomposition of the difference between population and empirical risk
$=$ $\langle\widehat{\mathcal{T}}_{K,0}(h_{0}-h+\eta),\widehat{\mathcal{T}}_{K,1}(h_{0}-h+\eta)\rangle-\langle\mathcal{T}_{K}(h_{0}-h),\mathcal{T}_{K}(h_{0}-h)\rangle$ (53)
$=$ $\langle(\widehat{\mathcal{T}}_{K,0}-\mathcal{T}_{K})(h-h_{0}),(\widehat{\mathcal{T}}_{K,1}-\mathcal{T}_{K})(h-h_{0})\rangle$ (54)
$+\sum_{v=0,1}\langle\mathcal{T}_{K}(h-h_{0}),(\widehat{\mathcal{T}}_{K,v}-\mathcal{T}_{K})(h-h_{0})\rangle$ (55)
$+\sum_{v=0,1}\langle\mathcal{T}_{K}(h-h_{0}),(\widehat{\mathcal{T}}_{K,v}-\mathcal{T}_{K})\eta\rangle$ (56)
$+\sum_{v=0,1}\langle(\widehat{\mathcal{T}}_{K,1-v}-\mathcal{T}_{K})(h-h_{0}),(\widehat{\mathcal{T}}_{K,v}-\mathcal{T}_{K})\eta\rangle$ (57)
$+\langle(\widehat{\mathcal{T}}_{K,0}-\mathcal{T}_{K})\eta,(\widehat{\mathcal{T}}_{K,1}-\mathcal{T}_{K})\eta\rangle$ (58)
@@ B.3 U-process terms
Denote $\mathcal{G}=\mathcal{H}-h_{0}$. For any $g\in\mathcal{G}$, let $g^{0}=g-\mathcal{T}_{K}g$, and let $\mathcal{G}^{0}=\{g-\mathcal{T}_{K}g:g\in\mathcal{G}\}$ For any such $g^{0}$, let
$Z_{N,1}^{0}(g)=\langle\widehat{\mathcal{T}}_{K,0}g^{0},\widehat{\mathcal{T}}_{K,1}g^{0}\rangle.$ (59)
@@
$[\widehat{\mathcal{T}}_{K,0}g^{0}](a)=O(\widetilde{\delta}_{n}\|g^{0}\|_{2,a}+\widetilde{\delta}_{n}^{2}).$ (62)
@@
$Z_{N,1}(g^{0})\leq\widehat{\delta}_{N}\|\widehat{\mathcal{T}}_{K,0}g^{0}\times g\|_{|\mathcal{D}_{0}}+\widehat{\delta}_{N}^{2}.$ (63)
@@
$\|\widehat{\mathcal{T}}_{K,0}g^{0}\times g\|_{|\mathcal{D}_{0}}=$ $\sqrt{\frac{2}{N}\sum_{i:V_{i}=1}\left\lvert\widehat{\mathcal{T}}_{K,0}g^{0}(A_{i})\right\rvert^{2}\mathbb{E}[(g^{0}(S_{i}))^{2}]}$ (64)
@@
$Z_{N,2}^{0}(g)=\langle\widehat{\mathcal{T}}_{K,0}\eta,\widehat{\mathcal{T}}_{K,1}g^{0}\rangle$ (68)
@@
$\max_{a\in[K]}|[\widehat{\mathcal{T}}_{K,0}\eta](a)|=O\left(\sqrt{\frac{\log K}{n}}\right).$ (70)
@@
$Z_{N,2}^{0}(g)=\frac{1}{N/2}\sum_{i:V_{i}=1}[\widehat{\mathcal{T}}_{K,0}\eta](A_{i})g^{0}(S_{i}).$ (71)
The conclusion then follows by applying lemma 9 conditional on $\mathcal{D}_{0}$, on the event that the above display holds, with $b_{i}=[\widehat{\mathcal{T}}_{K,0}\eta](A_{i})|$. ∎